OpenStudy (anonymous):

First on e answer and show work get a medal Divide 365 by 7 and write the anser as a mixed number. First on e answer and show work get a medal Divide 365 by 7 and write the anser as a mixed number. @Mathematics

jimthompson5910 (jim_thompson5910):

365/7 (364+1)/7 364/7+1/7 52+1/7 52 1/7 So \[\Large \frac{365}{7} = 52\frac{1}{7}\]